The jasthesis L^{A}TEX class

(How to typeset a thesis) Volume 1 of 1

James A. Shepherd

A dissertation submitted to the University of Life in accordance with the

requirements of… Word Count: many

Abstract

A L^{A}TEX class that conforms to the requirements of a University Thesis. It is approximately BS4821:1990.

Briefly these requirements are:

• A4 papersize

• Consecutive numbering at the bottom center of each page.

• Text in 1.5 line spacing (Titles and contents pages are still singlespaced)

• 12pt font size.

• Doublesided printing.

• Top, bottom and outer margin at least 15mm. Inner margin 40mm.

• Preliminary pages of:

– Title Page (listing chapters and sections)

– Abstract

– Dedication and acknowledgements

– Author’s Declaration

– Table of Contents, List of Figures, List of Tables

• Top left heading is the current chapter, top right current section

Note that when using the option jasheads the ∗-form of the section command is not supported properly.

Also, the requirement that headings be in no more than 14pt is ignored. This would look silly, given the large gaps in the text due to displayed mathematics. Though if you really want it, the class file would not take much editing to do it.

2

Dedication and Acknowledgements

This document class is dedicated to my girlfriend, Codina Cotar.

4

Author’s Declaration

I declare that the work in this dissertation was carried out in accordance with the Regulations of the…

Signed: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .
Date: . . . . . . . . . . . . . . . . . . . . .

6

Contents

1 Using the Class 9 1.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2 Title and Contents Pages 11 2.1 Title Page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.2 Contents Pages . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

3 Abstract, Dedication, and Declaration 13

4 Other 15 8

Chapter 1

Using the Class

Place a copy of the jasthesis.cls file in the same directory/folder as the .tex file you are editing. Alternatively, if you are able to, install the class in your L^{A}TEX source tree.

At the beginning of you L^{A}TEX use the command:

   \documentclass{jasthesis}
instead of your usual \documentclass{\ldots}.

1.1. Options

The option jasheads gives more modern headings on the top left and right. So, you would invoke this with:

   \documentclass{jasthesis}

instead of the above command.

All the remaining command will usually be used immediately after the above.

Chapter 1 Using the Class 10

Title and Contents Pages

2.1. Title Page

Use the following commands to enter the details for the title page:

\thesistitle{\ldots}
\thesissubtitle{\ldots}
\thesisvolume{\ldots}
\thesisauthor{\ldots}
\thesisdegree{\ldots}
\thesiswordcount{\ldots}

Then use the command \thesismaketitle to actually produce the page.

2.2. Contents Pages

Using the standard L^{A}TEX command \tableofcontents will produce the table of contents with chapters and sections. It will be only singlespaced.

If you require a single sided contents page, then this will have to be done manually. Where you want the page break on the contents page, go to that section in your .tex file and after the section command put the command:

\addtocontents{toc}{\cleardoublepage}

The standard L^{A}TEX commands \listoffigures and \listoftables may be used as usual.

Chapter 2 Title and Contents Pages 12

Abstract, Dedication, and Declaration

These details are entered using the commands:

\thesisabstract{\ldots}
\thesisdedication{\ldots}
\thesisdeclaration{\ldots}

These commands also produce the pages. The second one typesets the text in italics. The last one produces the signed and date lines.

Chapter 3 Abstract, Dedication, and Declaration 14

Chapter 4 Other

The rest of the text will be 1.5 linespaced. I guess that’s all.

http://www.jamesAshepherd.com/